The residual dividend policy approach to dividend policy is based on the theory that a firm’s optimal dividend distribution policy is a function of the firm’s target capital structure, the investment opportunities available to the firm, and the availability and cost of external capital. The firm makes distributions based on the residual earnings.

Consider the case of Red Bison Petroleum Producers Group:

Red Bison Petroleum Producers Group is expected to generate $140,000,000 in net income over the next year. Red Bison Petroleum Producers Group has forecasted a capital budget of $83,000,000, and it wishes to maintain its current capital structure of 70% debt and 30% equity. 30% Equity 70% Debt

If the company follows a strict residual dividend policy and makes distributions in the form of dividends, what is its expected dividend payout ratio for this year?

82.21%

86.32%

57.55%

73.99%

If Red Bison Petroleum Producers Group reduces the amount of its forecasted capital budget, how will this affect the firm’s annual dividend, assuming that all other factors are held constant?

The amount that Red Bison Petroleum Producers Group will pay out in dividends this year will decrease.

The amount that Red Bison Petroleum Producers Group will pay out in dividends this year will increase.

Most firms have earnings that vary considerably from year to year and do not grow at a reliably constant pace. Furthermore, their required investment may change often. Which of these statements is the most accurate?

Most firms can still use the concepts behind a residual dividend policy to make long-run decisions about dividends.

A residual dividend policy can’t be of any help to most firms.

Grade It Now

Save & Continue

Continue without saving
Answer #1

Under residual dividend policy, earnings remaining after meeting all profitable investments is distributed as dividend

Capital Budget = $83,000,000

To be financed through Equity = 83,000,000*30% = $24,900,000

Dividend = 140,000,000 – 24,900,000

= $115,100,000

Dividend payout ratio = 115,100,000/140,000,000

= 82.21%

Dividend will INCREASE

Most firms can still use the concepts behind a residual dividend policy to make long-run decisions about dividends.